# Quantum Rationale-Aware Graph Contrastive Learning for Jet Discrimination

### v6

@@ 3.3 Benchmark Models
[…]
4 Experimental Setup
In our experiments, we utilized various tools and frameworks to implement and train the models. The classical models were implemented using PyTorch 2.2.0 [27], while the quantum components were handled using Pennylane 0.38.0 [28], TorchQuantum 0.1.8 [29], and Qiskit [30]. To simulate graph-based operations, we used the Deep Graph Library (DGL) 2.1.0+cu121 [31]. The full experimental setup, including hardware specifications, hyperparameter configurations, and detailed experimental procedures, can be found in Appendix F.
Training protocol.
We adopt a two-stage procedure: (i) self-supervised pretraining for 50 epochs using a rationale-aware contrastive objective $\mathcal{L}_{\text{RA}}+\lambda\,\mathcal{L}_{\text{CP}}+\alpha\,\mathcal{L}_{\text{align}}+\beta\,\mathcal{L}_{\text{uniform}}+\delta\,\mathcal{L}_{\text{InfoNCE}}$, and (ii) supervised fine-tuning with a linear classifier for 1000 epochs on the learned graph-level embeddings (the encoder is frozen). This clarifies how representations are first aligned via CL and then evaluated for jet discrimination.
5 Results and Discussion
We evaluated the performance of QRGCL against two classical models (GNN and EQGNN), three quantum models (QGNN, EQGNN, and QCL), and five hybrid classical-quantum models (CQCL, three variants of QCGCL, and QRGCL with RX+H encoding). The area under the curve (AUC) was selected as the benchmark metric due to its effectiveness in assessing binary classification performance across all thresholds. The number of trainable parameters of CRG was 1,073, compared to 45 for QRG.
@@ 5 Results and Discussion
Figure 5: (a) Training and testing dynamics of QRGCL over 1000 epochs. The left graph illustrates loss curves, while the right graph presents accuracy curves. (b) ROC-AUC curve for all the benchmarked models
QRGCL achieved the highest AUC (77.53%), outperforming all baselines. The best performance was obtained using $H$ followed by $RX$ encoding, a learning rate of $1\times 10^{-3}$, and $SWAP$ entanglement, as detailed in Table 1. As shown in Table 2, QRGCL outperformed classical RGCL in most settings, particularly with 3 layers, 7-node rationales, and an augmentation ratio of 10%. More detailed results are outlined in Appendix G.
We took the best two well-performing encoders from Table 1 and developed two variants by hybridizing them, and further experimentations revealed that $H$ followed by $RX$ gate is more robust. Additionally, the relatively large number of parameters in the QRGCL is due to the utilization of the ParticleNet GNN encoder, which possesses 125k parameters, while the QRG circuit only has 45 parameters. Tables 1, 2, and 3 complement the ablation studies for QRGCL. Full ablation results, including encoder comparisons, entanglement strategies, layer and rationale variations, and rationale-aware vs. rationale-agnostic models, are detailed in Appendix G.
Additional benchmarks and resource notes (including cases where certain quantum baselines fail at larger $n_{\alpha}$ due to memory limits) are provided in Table 4 (and reproduced in Appendix A).
@@ 7 Conclusion
This paper introduced QRGCL, a hybrid quantum-classical rationale-aware GNN for quark-gluon jet classification, integrating quantum computation with GCL. Our results show that QRGCL outperforms two classical, three quantum, and five hybrid classical-quantum models, achieving an AUC of 77.53%. Hyperparameter tuning indicated that the $H+RX$ encoder and the $SWAP$ entanglement gate optimized performance, emphasizing the importance of careful circuit design in quantum models. QRGCL performed best on smaller graphs (7 nodes) with low augmentation (0.1), achieving an AUC of 78.78%, indicating their value for compact, high-quality representations. Ablation studies showed that the QRG benefits from a 3-layer architecture, while the CRG performs best with four layers, underscoring the potential of quantum-enhanced approaches in particle physics.
@@ 8 Limitations
Future work could focus on reducing parameter complexity to streamline model efficiency, try other state-of-the-art encoder networks like LorentzNet [32] and Lorentz-EQGNN [33], and also the hybridization of RG. We also plan to extend the model to other high-energy physics tasks, such as anomaly detection in particle collisions and event reconstruction. We look forward to improving the explainability of GCL and exploring how retrospective and introspective learning in rationale discovery can guide discrimination tasks and enhance the generalization of backbone models.
Appendix A Additional Benchmark Results
Table 4: Performance comparison of the benchmarked models for $n_{\alpha}>3$.
Model Test Acc AUC F1 score $n_{\alpha}$ #Parameters $n_{\text{layer}}$ Batch size Epoch Encoder
EGNN 54.2% 64.39% 68.52% 7 5252 5 64 19 MLP
EQGNN 47.8% 55.13% 30.92% 7 990100 6 1 19 MLP + H
QGNN 54.6% 43.90% 54.62% 7 1021076 6 1 19 MLP + H
GNN 52.2% 57.76% 35.81% 7 5252 4 64 19 MLP
QCL 44.8% 48.28% 61.88% 5 384 3 256 1000 Amplitude
QCL 7 3 256 1000 Amplitude
CQCL 45.2% 48.42% 60.74% 5 354 3 128 50 Angle (RY+RX)
CQCL 7 3 128 50 Angle (RY)

### v7

@@ 3.3 Benchmark Models
4 Experimental Setup
4.1 Simulation Tools and Environment
We implemented all the models using the PyTorch 2.2.0 (pytorch) framework for classical computations and Pennylane 0.38.0 (bergholm2018pennylane) and TorchQuantum 0.1.8 (torchquantum) for quantum circuit simulation. We used the Deep Graph Library (DGL) 2.1.0+cu121 (DGL) to handle graph operations and the Qiskit 0.46.0 (qiskit2024) framework to simulate quantum circuits. The computing infrastructure consisted of Intel(R) Xeon(R) CPUs (x86) with a clock frequency of 2 GHz, equipped with 4 vCPU cores and 30 GB of DDR4 RAM. For GPU acceleration, we utilized two NVIDIA T4 GPUs, each with 2560 CUDA cores and 16 GB of VRAM, significantly boosting the performance of deep learning tasks. For reproducibility, we set random seeds to 42, 123, and 456 for the 3-seed averaging reported in Table 1. Common training utilities included: (i) learning rate warm-up with linear increase from $1\times 10^{-6}$ to base LR over specified epochs and (ii) gradient clipping. Quantum circuit differentiation exclusively utilized the parameter-shift rule to ensure hardware compatibility, with statevector simulation employed for models requiring exact gradients and shot-based simulation (with 1024 shots) specified. The source code for QRGCL, including the quantum rationale generator implementation and benchmarking suites, is publicly available at https://github.com/Abrar2652/QRGCL.
4.2 Hyperparameters and Configurations
We varied hyperparameters for CRG, QRG, data augmentation, and training in the proposed QRGCL model, and reported test accuracy, AUC, and F1 scores across various encoder types, learning rates (LR), and entanglement strategies. We used Adam optimizer with a $1\times 10^{-3}$ learning rate ($\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$) across all the models and a Binary Cross-Entropy (BCE) loss function. 10-fold cross-validation was performed for each model, and the mean and standard deviation were calculated for each metric. The hidden feature size for classical GNN and EGNN was 10 to maintain a comparable parameter count to the quantum counterparts, while for the QGNN and EQGNN, it was $2^{n_{\alpha}}$=8. In the case of QCL and CQCL, $n_{layer}$ stands for the depth of the quantum circuit, and in other models, it refers to the number of GNN layers. The term $n_{layer}$ in QRG refers to the number of repetitions of the variational block, which consists of the parameterized rotation gates ($U3$) and the entanglement layers. A deeper circuit ($n_{layer}>1$) increases the expressivity of the quantum ansatz. In Table 2, each row varies a single hyperparameter while keeping all others fixed to a shared default configuration ($n_{\alpha}$ = 7 nodes, two GNN layers, 10% dropout, GAT encoder, embedding dimension 128, hidden size 256, output dimension 128, no weight decay, augmentation ratio 0.1, temperature 0.1, batch size 2000, and for QRGCL only: 3 quantum layers). Specifically, the third row sets the number of GNN layers to 4 for CRGCL, while all other parameters remain at their default values. No joint (pairwise) hyperparameter optimization was performed due to computational constraints; thus, comparisons reflect controlled one-factor-at-a-time sweeps under consistent training settings.
The choice of epochs and batch sizes varies based on the computational requirements of classical and quantum models. Classical models like GNN and EGNN use larger batch sizes (64) and fewer epochs (19) to achieve more efficient gradient updates, enabling faster convergence. In contrast, quantum models such as QGNN and EQGNN use much smaller batch sizes (1) due to quantum hardware limitations, yet still require 19 epochs to achieve adequate learning despite fewer updates per step. For fully quantum and hybrid models like QRGCL and QCGCL, larger batch sizes (2000) and more epochs (50) are used due to the longer backpropagation caused by the complex custom loss function, with multiple loss components ensuring that the model has sufficient time to learn robust quantum-based representations, as the slower convergence can hinder effective learning.
Training Protocol.
For fair comparison, classical GNN, EGNN, and EQGNN baselines were trained using fully supervised learning, whereas QCL, CQCL, and QCGCL adopted contrastive objectives. Only QRGCL and its classical counterpart CRGCL employed rationale-aware contrastive pretraining followed by fine-tuning, isolating the contribution of the QRG. All supervised benchmarks were trained using BCE loss with the Adam optimizer ($lr=10^{-3}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$). Contrastive benchmarks (QCL, CQCL) utilized InfoNCE loss ($T=0.1$) during pre-training. We adopt a two-stage procedure for QRGCL: (i) self-supervised pretraining for 50 epochs using a rationale-aware contrastive objective defined as $\mathcal{L}_{\text{QRGCL}}\ =(\mathcal{L}_{\text{RA}}+\lambda\,\mathcal{L}_{\text{CP}}+\alpha\,\mathcal{L}_{\text{align}}+\beta\,\mathcal{L}_{\text{uniform}}+\delta\,\mathcal{L}_{\text{InfoNCE}}$), and (ii) supervised fine-tuning with a linear classifier for 1000 epochs on the learned graph-level embeddings (during fine-tuning, the encoder parameters are kept fixed). Convergence of the contrastive loss was monitored using a validation set during the 50-epoch pretraining phase to ensure sufficient feature alignment before the extensive 1000-epoch fine-tuning stage. This step evaluates the quality of learned representations in a supervised downstream setting. Figure 5 illustrates the learning dynamics over 1000 epochs. While the accuracy begins to stabilize around 500–800 epochs, we extended training to 1000 epochs to ensure complete convergence of the projection head. Crucially, no degradation in validation performance was observed in the later epochs, indicating that the model is robust against overfitting even with prolonged training.
Scalability.
Models ending with “CL” in Table 3 used CL but lacked rationale generation. For other models, increasing $n_{\alpha}$ corresponds to including more raw particle features in the encoder without explicitly learning which ones are most informative. As $n_{\alpha}$ increases, both quantum and hybrid models experience exponential scaling in circuit width and memory usage, making larger graphs harder to simulate reliably and resulting in failed runs for QCL and CQCL at $n_{\alpha}=7$. It is important to note that this exponential bottleneck applies to the classical simulation of the quantum state. On actual quantum hardware, the resource scaling would be linear with respect to the number of nodes (qubits), limited primarily by gate fidelity and coherence times rather than memory. In contrast, classical graph-based models such as GNN and EGNN maintain stable accuracy and F1-scores, reflecting linear computational scaling. QGNN and EQGNN remain trainable but incur a steep parameter cost ($\approx 10^{6}$) without proportional performance gains. The absence of consistent improvement, and occasional degradation, with larger $n_{\alpha}$ supports our central finding: simply enlarging the particle subgraph does not improve jet discrimination unless the most salient constituents are identified, as done by QRGCL. The observation that performance does not improve and often worsens with higher $n_{\alpha}$ supports our main claim: larger input subgraphs do not necessarily improve discrimination in jet data unless the substructure is meaningfully selected, as in QRGCL.
5 Results and Discussion
We evaluate the performance of QRGCL against the baseline architectures defined in Section 3.3 (see also Appendix B.3 and C for detailed specifications). AUC was selected as the benchmark metric due to its effectiveness in assessing binary classification performance across all thresholds. While specific working points (e.g., background rejection at fixed efficiency) are often used in HEP, AUC provides a holistic measure of the discriminator’s separation power independent of specific operating conditions, facilitating architectural comparison. The number of trainable parameters of CRG was 1,073, compared to 45 for QRG.
Table 4: Hyperparameter optimization results of the proposed QRGCL. Reported values represent the mean validation accuracy across stratified 10-fold cross-validation (with shuffling enabled) on the training set. Standard deviations indicate the variability across folds. Bolded values indicate the best performance. Unless otherwise varied in a specific row, the default configuration uses $n_{\alpha}=7$ nodes, 3 quantum layers, a learning rate of $1\times 10^{-3}$, SWAPSWAP entanglement, and RXRX encoding.
Parameter type Parameter Accuracy ($\uparrow$) AUC ($\uparrow$) F1 Score ($\uparrow$)
Encoder type Amplitude $69.3\%\pm 1.1\%$ $74.7\%\pm 1.0\%$ $69.3\%\pm 1.1\%$
IQP $68.9\%\pm 1.2\%$ $76.1\%\pm 0.9\%$ $68.5\%\pm 1.2\%$
Displacement-amplitude $66.2\%\pm 1.5\%$ $72.1\%\pm 1.4\%$ $66.0\%\pm 1.6\%$
Displacement-phase $70.2\%\pm 1.0\%$ $76.2\%\pm 0.9\%$ $69.1\%\pm 1.0\%$
RY $67.0\%\pm 1.3\%$ $73.9\%\pm 1.1\%$ $67.0\%\pm 1.2\%$
RX $\mathbf{70.8\%\pm 0.9\%}$ $\mathbf{76.3\%\pm 0.8\%}$ $\mathbf{69.6\%\pm 0.8\%}$
RZ $62.9\%\pm 2.1\%$ $67.3\%\pm 1.9\%$ $62.8\%\pm 2.0\%$
H $\mathbf{70.8\%\pm 0.8\%}$ $\mathbf{76.7\%\pm 0.7\%}$ $\mathbf{69.7\%\pm 0.8\%}$
Phase $68.5\%\pm 1.2\%$ $74.2\%\pm 1.1\%$ $68.3\%\pm 1.2\%$
Learning rate 1E-04 $68.6\%\pm 1.1\%$ $74.4\%\pm 1.0\%$ $68.4\%\pm 1.1\%$
1E-03 $\mathbf{70.3\%\pm 0.8\%}$ $\mathbf{76.2\%\pm 0.8\%}$ $\mathbf{70.3\%\pm 0.8\%}$
3E-03 $63.9\%\pm 1.6\%$ $68.8\%\pm 1.5\%$ $63.7\%\pm 1.7\%$
5E-03 $66.0\%\pm 1.4\%$ $70.9\%\pm 1.3\%$ $66.0\%\pm 1.4\%$
1E-02 $69.5\%\pm 1.0\%$ $75.8\%\pm 0.9\%$ $69.5\%\pm 1.0\%$
Entanglement type CNOT $69.8\%\pm 1.0\%$ $76.6\%\pm 0.9\%$ $69.8\%\pm 1.0\%$
CZ $68.2\%\pm 1.2\%$ $73.5\%\pm 1.1\%$ $68.1\%\pm 1.2\%$
SWAP $\mathbf{71.0\%\pm 0.9\%}$ $\mathbf{76.2\%\pm 0.8\%}$ $\mathbf{70.8\%\pm 0.9\%}$
CNOT Butterfly $68.0\%\pm 1.3\%$ $74.6\%\pm 1.2\%$ $68.0\%\pm 1.3\%$
CZ Butterfly $67.7\%\pm 1.4\%$ $73.2\%\pm 1.3\%$ $67.7\%\pm 1.4\%$
SWAP Butterfly $68.0\%\pm 1.3\%$ $73.2\%\pm 1.2\%$ $67.8\%\pm 1.3\%$
@@ 5 Results and Discussion
Figure 5: Training and testing dynamics of QRGCL over 1000 epochs. The left graph illustrates loss curves, while the right graph presents accuracy curves.
Figure 6: ROC curves comparing the proposed QRGCL with classical, quantum, and hybrid baseline models on the quark–gluon jet discrimination task. QRGCL achieves competitive AUC performance across a wide range of operating points. In particular, QRGCL maintains a higher TPR at low FPRs, highlighting its effectiveness in regimes relevant for high-purity jet tagging. These results confirm that incorporating quantum-enhanced rationale selection can improve discriminative performance while remaining parameter-efficient.
The hyperparameters specific to QRGCL include encoder type, learning rate, and entanglement type, as detailed in Table 4. See Appendix B.3 for detailed definitions of the specific quantum encoders tested, including Amplitude, IQP, and Angle-based variants. Among the encoders, $RX$ and $H$ achieved the highest AUC values at 76.3% and 76.7%, respectively, while displacement-amplitude and $RZ$ encodings performed poorly, with AUCs of 72.1% and 67.3%. The optimal learning rate of $1\times 10^{-3}$ produced the highest AUC of 76.2%, with higher rates resulting in decreased performance. The $SWAP$ entanglement type yielded the best overall results, achieving an AUC of 76.2%. $CNOT$ and $CZ$ entanglements performed strongly, while other configurations underperformed. For the proposed QGRCL, two possible combinations of $RX$ and $H$ encoders, the learning rate of $1\times 10^{-3}$ and $SWAP$ entangler were tried.
For both the classical RGCL and QRGCL models, tunable parameters included the number of nodes per head ($n_{\alpha}$) ranging from 7 to 10, the number of classical or quantum layers ($n_{layer}$) ranging from 2 to 5, and the augmentation ratio ranging from 0.1 to 0.3. As shown in Table 2, QRGCL achieves comparable AUC values to the classical RGCL across different node counts (8 and 10 nodes). At 3 layers, QRGCL achieves performance comparable to the classical model (74.7% vs 73.6%, within statistical uncertainty). However, it underperformed at 2 and 4 layers. With a 0.1 augmentation ratio, QRGCL achieved its strongest performance of 78.8%, compared to the classical approach’s 74.1%. The optimal parameters for QRGCL were found to be $n_{\alpha}=7$, $n_{layer}=3$, and an augmentation ratio of 10%.
Based on Table 1 and Figure 6, it is evident that our proposed QRGCL model achieves a robust mean AUC score of 77.5% (averaged over 3 seeds), with peak performance reaching 78.8% in optimal configuration runs (Table 2). While QRGCL achieves the highest mean AUC, the error margins indicate competitive performance with state-of-the-art classical models, suggesting that the quantum advantage lies primarily in parameter efficiency and rationale interpretability rather than raw accuracy alone. We selected the two best-performing encoders from Table 4 and developed two variants by hybridizing them. Further experimentation revealed that $H$ initialization followed by $R_{X}$ encoding demonstrated greater stability, achieving the highest mean AUC and lowest standard deviation across validation folds. Additionally, the relatively large number of parameters in the QRGCL is due to the utilization of the ParticleNet GNN encoder, which contains 125k parameters, while the QRG circuit only has 45 parameters. Tables 1,2, and 4 complement the ablation studies for QRGCL by presenting results for configurations without rationale-awareness (QGNN, GNN, EGNN, EQGNN), analyzing VQC components (such as encoding variants, entanglement structures, and variations in qubits and layers), exploring different classical-quantum interfaces (including quantum-only and hybrid architectures, as well as a classical-only baseline), and examining the effects of rationale-guided data augmentation. Figure 5 shows steadily decreasing training and testing losses in QRGCL, indicating effective learning. While the training and testing curves stabilize around 800 epochs, we observe a persistent gap ($\sim 2\%$) and a slight decline in test accuracy in later epochs, indicating mild overfitting typical of deep learning models trained on restricted datasets (10k samples). Additionally, we observe that classical GNNs perform slightly better in the high-purity regime (tight cuts, Figure 6 bottom-left), likely because they utilize the full feature set without the information bottleneck imposed by the quantum rationale generator. Training and testing accuracies increase rapidly within the first 200 epochs and plateau near 800 epochs. Test accuracy stabilizes slightly above 70%, with training accuracy close behind, indicating stable performance and limited gains from further training. Figure 6 presents the ROC curves for the top-performing models, highlighting that QRGCL (blue curve) maintains a competitive true positive rate across most false positive thresholds compared to the classical GNN and hybrid baselines.
@@ 7 Conclusion
This paper introduced QRGCL, a resource-constrained rationale-aware graph contrastive learning framework that integrates a quantum rationale generator into a hybrid quantum–classical pipeline. We evaluate this methodology on quark–gluon jet discrimination as a representative use case, demonstrating that rationale-aware contrastive pretraining with a compact QRG can yield competitive downstream performance. Our results show that QRGCL achieves a robust mean AUC of 77.5%, with peak performance reaching 78.8% in optimal configuration runs, while keeping the quantum rationale module extremely lightweight (45 parameters). Hyperparameter analysis further indicates that the H+RXH{+}RX encoder and the $SWAP$ entanglement gate improve stability and performance, emphasizing the importance of circuit design choices in this regime. These findings support QRGCL as a general approach for learning representations from graph-structured data when resources (parameters, simulation cost, or supervision) are limited, with jet tagging serving as a concrete and challenging testbed.
@@ 8 Limitations
A primary limitation of this study is the use of generator-level data and aggressive jet grooming ($n_{\alpha}=7$) necessitated by current quantum simulation constraints. While real LHC analyses involve full detector reconstruction and $O(50)$ particles, this work serves as a proof-of-concept for the parameter efficiency of Quantum Rationales. Our experiments are limited to a narrow $p_{T}$ range (500-550 GeV) and exclude detector effects/pileup, which limits conclusions regarding real-world LHC performance. The use of only 10k training samples, necessitated by quantum simulation costs, may disadvantage the larger classical benchmark models. Additionally, our current QRG architecture utilizes a classical linear layer to project the 8-dimensional particle feature vectors onto a single scalar for angle encoding. While this ensures parameter efficiency, it acts as an information bottleneck that may limit the quantum circuit’s ability to leverage complex nonlinear correlations between the original raw features compared to a full-width classical encoder. Future work could focus on reducing parameter complexity to streamline model efficiency, exploring other state-of-the-art encoder networks, such as LorentzNet (gong_efficient_2022) and Lorentz-EQGNN (Jahin_2025), and also hybridizing RG. We also plan to extend the model to other high-energy physics tasks, such as anomaly detection in particle collisions and event reconstruction. We look forward to improving the explainability of GCL and exploring how retrospective and introspective learning in rationale discovery can guide discrimination tasks and improve the generalization of backbone models.
Appendix A Details of Dataset and Preprocessing
A.1 Details of High-Energy Physics Dataset
This study uses the Pythia8 Quark and Gluon Jets for Energy Flow (komiske_energy_2019) dataset, a well-established benchmark in high-energy physics for jet classification. The dataset comprises 2 million simulated particle jets, evenly divided into 1 million quark-originated jets and 1 million gluon-originated jets. The dataset was generated using Pythia 8.226 with the default Monash 2013 tune. The hard-scatter processes simulated were $Z(\to\nu\bar{\nu})+(u,d,s)$ for quark jets and $Z(\to\nu\bar{\nu})+g$ for gluon jets, computed at leading order (LO). Hadronization was performed using the Lund string model. The quark jets consist exclusively of light quarks ($u,d,s$) and do not include heavy flavor ($c,b$) jets. No detector simulation or pileup interactions were included. These jets are generated through collision events simulating the Large Hadron Collider (LHC) conditions with a center-of-mass energy $\sqrt{s}=14~\text{TeV}$. Jets were selected based on their transverse momentum range $p_{T}^{\text{jet}}$ between 500 and 550 GeV and their pseudorapidity $|y^{\text{jet}}|<1.7$. Each jet $\alpha$ is labeled as a quark jet ($y_{\alpha}=1$) or a gluon jet ($y_{\alpha}=0$), providing a binary classification target for model training.
Each particle $i$ within a jet is characterized by several key attributes: transverse momentum $p_{T,\alpha}^{(i)}$, rapidity $y_{\alpha}^{(i)}$, azimuthal angle $\phi_{\alpha}^{(i)}$, and its Particle Data Group (PDG) identifier $I_{\alpha}^{(i)}$. The particle mass $m_{\alpha}^{(i)}$ is derived directly from the PDG identifier, allowing for the complete reconstruction of the 4-momentum vector used in subsequent feature engineering (Eq. 7). From these basic features, we derive additional physically motivated quantities including particle masses (from PDG identifiers), transverse masses $m_{T,\alpha}^{(i)}=\sqrt{m_{\alpha}^{(i)2}+p_{T,\alpha}^{(i)2}}$, energies $E_{\alpha}^{(i)}=m_{T,\alpha}^{(i)}\cosh y_{\alpha}^{(i)}$, and momentum components.
Figure 7: Aggregate kinematic distributions of the Pythia-generated dataset: (a) Particle multiplicity across all jets; (b) Transverse momentum ($p_{T,\alpha}$); (c) Total momenta ($p_{\alpha}$); and (d) Energy ($E_{\alpha}$). These distributions represent the combined quark and gluon populations before grooming and class-specific analysis.
A.1.1 Comparative Analysis of Quark and Gluon Jets
General kinematic distributions for the entire simulated dataset, including particle counts and energy ranges before class separation, are visualized in Figure 7. To provide a comprehensive understanding of the underlying class structure and the distinguishing features between quark and gluon jets, we present detailed comparative analyses in Figures 1, 8, and 9.
Figure 1 presents the four most discriminating features that clearly distinguish quark and gluon jets:
  • Particle Multiplicity (Fig. 1a): Gluon jets contain approximately 59% more particles on average than quark jets ($\mu_{\text{gluon}}=53.2$ vs. $\mu_{\text{quark}}=33.4$). This fundamental difference arises from the QCD color factors: gluons carry a color charge of 8 (adjoint representation with $C_{A}=3$) compared to quarks’ charge of 3 (fundamental representation with $C_{F}=4/3$). While the color factor ratio $C_{A}/C_{F}=9/4\approx 2.25$ predicts that gluons radiate approximately 2.25 times more strongly than quarks, the observed 59% difference in this study is attributed to hadronization effects and the $n_{\alpha}=7$ truncation (grooming) required for quantum simulation. This truncation specifically suppresses the soft radiation contributions that typically drive higher multiplicity in gluon jets.
  • Jet Width (Fig. 1b): Gluon jets exhibit a 46% larger spatial extent in the $\eta$-$\phi$ plane compared to quark jets ($\mu_{\text{gluon}}=0.089$ vs. $\mu_{\text{quark}}=0.061$). This increased width reflects the stronger coupling of gluons to the color field and their enhanced emission of soft, wide-angle radiation. The jet width is computed as $w=\sqrt{\langle\Delta y^{2}+\Delta\phi^{2}\rangle}$, where the angular distances are weighted by particle transverse momentum.
  • Leading Particle $p_{T}$ Fraction (Fig. 1c): Quark jets demonstrate significantly harder fragmentation, with the leading particle carrying 63% more of the total jet momentum on average than in gluon jets ($\mu_{\text{quark}}=0.284$ vs. $\mu_{\text{gluon}}=0.174$). This reflects the fact that quarks tend to fragment into a dominant hadron that preserves much of the original quark’s momentum, while gluons distribute their energy more democratically among multiple softer particles through cascading radiation.
  • Jet Mass (Fig. 1d): Gluon jets have 40% larger invariant mass compared to quark jets ($\mu_{\text{gluon}}=52.8$ GeV vs. $\mu_{\text{quark}}=37.6$ GeV). This increased mass is consistent with their higher multiplicity and softer particle spectra, as more particles with broader angular distribution contribute to larger invariant mass through $m^{2}=E^{2}-\vec{p}^{2}$.
Figure 8: Comprehensive comparison of twelve features characterizing quark and gluon jets. Each panel shows normalized distributions (probability density) for both classes, with statistical significance indicated by Kolmogorov-Smirnov test p-values. Top row: fundamental jet properties including multiplicity, mass, transverse momentum, and spatial width. Middle row: particle-level $p_{T}$ characteristics showing mean, standard deviation, leading particle value, and leading particle fraction. Bottom row: fragmentation and energy properties including top-3 particle fraction and energy distributions. All features show highly significant differences between classes ($p<10^{-10}$), demonstrating the rich discriminative structure in the dataset. Quark jets (red) consistently show harder, more collimated fragmentation patterns, while gluon jets (teal) exhibit softer, more diffuse radiation.
Figure 8 provides a systematic overview of all twelve engineered features across both jet classes. Only the final 8 particle-level features are used as model inputs; the remaining variables are shown for completeness and comparison. Beyond the four key features discussed above, several additional discriminating characteristics are evident:
  • Particle $p_{T}$ Statistics: The mean particle $p_{T}$ is lower in gluon jets (reflecting their higher multiplicity and softer spectrum), while the standard deviation is higher (indicating greater variation in particle energies). The leading particle $p_{T}$ shows clear separation, with quark jets having significantly higher values.
  • Fragmentation Patterns: The top-3 particle $p_{T}$ fraction demonstrates that not only the leading particle, but the entire high-$p_{T}$ component is more prominent in quark jets. This indicates fundamentally different fragmentation dynamics between the two parton types.
  • Energy Distributions: Mean particle energies show similar patterns to $p_{T}$ distributions, with gluons having lower mean values but higher standard deviations. Total jet energy, constrained by the selection criteria ($p_{T}^{\text{jet}}\in[500,550]$ GeV), shows relatively similar distributions but with gluons extending to slightly higher values due to their larger mass contributions.
All features demonstrate highly significant statistical differences between classes (Kolmogorov-Smirnov test $p<10^{-10}$), confirming that the dataset contains rich discriminative structure across multiple complementary observables.
Figure 9: Two-dimensional feature space projections revealing class separation patterns. Each panel shows scatter plots of 5,000 randomly sampled jets from each class. (a) Multiplicity vs. width shows clear separation with gluons occupying the high-multiplicity, large-width region. (b) Multiplicity vs. mass demonstrates correlated increases in both quantities for gluons. (c) Fragmentation (leading $p_{T}$ fraction) vs. width shows anti-correlated behavior, with harder fragmentation (higher fraction) corresponding to narrower jets. (d) Width vs. mass shows strong positive correlation, particularly for gluons. (e) Mean $p_{T}$ vs. energy reveals the kinematic relationships between transverse and total energy. (f) Multiplicity vs. fragmentation shows clear class separation, with quarks clustering at high fragmentation fractions and low multiplicities. The scatter patterns demonstrate that multivariate combinations of features provide improved discriminative power compared to single features alone.
Figure 9 explores the multivariate structure of the feature space through two-dimensional projections. These visualizations reveal several important patterns:
  1. 1.
    Correlated Discrimination: Multiple feature pairs show clear class separation, indicating that the discriminative information is distributed across many complementary dimensions. For example, the multiplicity-width plane (Fig. 9a) shows that gluons consistently occupy the high-multiplicity, large-width quadrant, while quarks cluster in the low-multiplicity, narrow-jet region.
  2. 2.
    Physical Correlations: The strong correlation between jet width and mass (Fig. 9d) reflects the physical connection between spatial extent and invariant mass, broader jets naturally have larger masses due to the geometric contribution to $m^{2}$. Similarly, multiplicity and mass are correlated (Fig. 9b) as more particles generally contribute to larger total mass.
  3. 3.
    Anti-correlations: The negative correlation between fragmentation hardness and jet width (Fig. 9c) demonstrates that jets with more concentrated energy (high leading $p_{T}$ fraction) tend to be more collimated (narrow), consistent with the physical picture of hard fragmentation producing tightly-clustered particle showers.
  4. 4.
    Class Overlap: While clear separation exists in feature space, there is non-trivial overlap between the classes, particularly in the intermediate regions of the parameter space. This overlap presents a genuine classification challenge and explains why sophisticated machine learning approaches are necessary to achieve optimal discrimination.
A.1.2 Statistical Significance and Effect Sizes
To quantify the discriminative power of each feature, we computed Cohen’s $d$ effect sizes and performed two-sample Kolmogorov-Smirnov (KS) tests. Table 5 summarizes these statistics for all features. The results show that:
  • All features exhibit highly significant differences (KS test $p<10^{-100}$), far exceeding conventional significance thresholds.
  • Particle multiplicity shows the largest effect size ($d=1.89$), indicating that this single feature provides strong discriminative power.
  • Multiple features demonstrate large effect sizes ($|d|>0.8$), suggesting that an optimal classifier should leverage the information from multiple complementary observables.
  • The relative differences between classes range from 20% to 63%, with the most dramatic differences appearing in multiplicity (+59.4%) and leading $p_{T}$ fraction (+63.2%).
Table 5: Statistical comparison of quark and gluon jet features. Cohen’s $d$ measures effect size (standardized difference between means), with $|d|>0.8$ indicating large effects. All features show highly significant differences ($p<10^{-100}$). Relative differences are computed as $\frac{100\times(\mu_{\text{gluon}}-\mu_{\text{quark}})}{\mu_{\text{quark}}}$.
Feature Quarks (mean $\pm$ std) Gluons (mean $\pm$ std) Relative Diff (%) Cohen’s $d$
Multiplicity $33.4\pm 10.6$ $53.2\pm 12.8$ +59.4 1.89
Jet $p_{T}$ [GeV] $524.2\pm 14.4$ $524.9\pm 14.5$ +0.1 0.05
Jet Mass [GeV] $37.6\pm 15.9$ $52.8\pm 18.3$ +40.4 0.98
Jet Width $061\pm 0.028 0.0$$089\pm 0.034$ +45.9 1.01
Mean Particle $p_{T}$ [GeV] $15.8\pm 7.4$ $10.1\pm 3.8$ -36.1 -1.12
Leading $p_{T}$ [GeV] $172.3\pm 89.4$ $92.4\pm 41.8$ -46.4 -1.28
Leading $p_{T}$ Fraction $0.284\pm 0.134$ $0.174\pm 0.074$ -38.7 -1.11
Top-3 $p_{T}$ Fraction $0.523\pm 0.168$ $0.365\pm 0.110$ -30.2 -1.21
Total Jet Energy [GeV] $759.2\pm 149.8$ $821.4\pm 171.3$ +8.2 0.40